The dawn reported (jan 17, 2022) that motorway ridership declined after a toll increases.



" There were nearly four million less riders in december 2021, the first full month after the price of a toll tax was increased 25% to Rs 125/10km, than in the previous month November 2021 in which nearly 40 million riders were recorded by the motorway Authority"



a) use these data to estimate the price elasticity of demand for motorway riders.



b) According to your estimate, what happens to the NHA's (National Highway Authority) revenue when the toll tax rises, graphically show the total revenue and give your opinion about authority decision?



c) do you think that your estimate of the elasticity is reliable?

1
Expert's answer
2022-02-10T13:57:36-0500

a) Price elasticity of demand


PED = The percentage change in quantity demanded divided by the percentage change in price 

40 - 4 = 36

percentage change in quantity demanded"= \frac{36}{(40+36)/2}" "\times100"

= -10.53%


Percentage change in Price ="\frac{2.5}{11.25}\times100"


PED= "\frac{-10.53}{22.22}"


PED= -0.47%


Decrease of 0.47



b)





The total revenue declined drastically. This is is because the the demand for motorway ridership decreased due to the increase in the price. Therefore, the government should reconsider adjusting the price for motorway ridership downward in order to attract more customers.


c) No. This is because the price level of an item affects the demand for a good or service, and the price elasticity of demand can be used to measure the sensitivity of a change in the quantity demanded of a good or service relative to a change in price. Therefore, the elasticity was not reliable.